# Born a Transformer – Always a Transformer?

### v2

@@ C.3 Promping: Lorem Ipsum Copying
To stress-test copying while keeping text realistic, we construct a 1,500-example corpus of synthetic Lorem Ipsum paragraphs. Generation starts from one base paragraph produced by the lorem Python library and then applies light, probabilistic perturbations to increase the non-determinism of the texts.
@@
  • Token budget. Sequences are truncated to 500 BPE tokens to stay within context limits while preserving paragraph-level coherence.
@@
We identified successful prompts for copying naturalistic text as follows. We downloaded research papers put up on ArXiv in April 2025, and checked if the prompt results in perfect accuracy (not even a single mistake). Our motivation was that ArXiv text incudes challenging text with equations, and, due to recency, should not have appeared in the LLM training data. More specifically, we segregated the sections of the papers into paragraphs, making sure the total length of the paragraph to be copied has close to 500 tokens (while ensuring that we do not end the paragraph in the middle of a sentence). We then pre-validated our prompts on a dataset containing $500$ samples each containing $500$ tokens, and even the smaller models we test – Llama3.1-8B and Qwen2.5-7B perfectly copied the given text (not a single mistake). We proceeded with our 3 minimal prompts, shown in Table 8, which achieved perfect copying accuracy. Each of our prompts ends with an unresolved <start> token, prompting the model to emit a verbatim repetition of the preceding paragraph.
Ratio of Ambiguous vs. Unambiguous Tokens. Table 7 reports, for each model, the average number of ambiguous (tokens with multiple possible continuations) and unambiguous (tokens with a single continuation) tokens per paragraph, aggregated across all random seeds. While different model families employ distinct tokenizers—so their absolute counts vary—the relative balance between ambiguous and unambiguous tokens is remarkably consistent. Each paragraph contains roughly 500 tokens, but repetition of sentences (to make the dataset appropriate for Non-unique copying) reduces the number of unique token types to only about 70. However, amongst those, both token types occur frequently in our data—thus, the fact that copying “glitch” can be traced exclusively to unambiguous tokens is remarkable.
@@
Table 7: Average per-paragraph bigram counts by model
@@
echo-tag RULE : Paragraphs are repeated **exactly** as is, whenever an Echo is seen.
@@
p-tag-the_end Please repeat the following paragraph exactly as it is, without any changes:
@@
C.4 Prompting: Git Commit Revert vs Cherry-Pick
@@
We synthesise commit logs by creating a randomized seven-character hexadecimal hash, an action/topic message drawn from 12 verbs and 10 subsystems. A history snippet lists $\ell$ commits newest,$\to$ oldest. We export three depths—$\ell=10,15,20,25$—and three random seeds, yielding $4\times 3\times 1,500=18,000$ total examples. For each snippet consisting of this generated git commit history (listed newest $\to$ oldest ) we store two reference answers:
  • revert: the lines as shown (newest,$\to$,oldest), matching the order a developer would pass to git revert (thus somewhat similar to forward copying)
  • cherrypick: the same lines reversed (oldest,$\to$,newest), the order expected by git cherry-pick. (thus somewhat similar to backwards copying)
@@
We adopt exactly two natural- language templates—one for the cherry-pick orientation, one for the revert orientation. Each prompt: describes the task in plain English, includes up to three few-shot examples (hand-crafted, $\ell=6$), shows the new history under a $=$ NEW HISTORY $=$ delimiter, and terminates with $=$ ANSWER $=$, where the model must emit the ordered commit list only. An example is given below.
@@
Because correct output demands global re-ordering rather than token-level copy, we found no benefit in adding separator toggles or mathematical re-statements.
@@
Appendix D Fine-Tuning Details
@@ D.4 Results
Figure 13: Averaged training loss curve for GPT-2-XL model across all our tasks. For all the tasks, we achieve 0 training loss. In contrast, generalization behavior varies in line with theoretical predictions (Figure 6).
Appendix E Further Details on Mechanistic Interpretability Analysis
@@
To remove an induction head, we use path patching methodology, widely adopted in the literature [Hanna et al., 2023, Wang et al., 2022]. An induction head can be viewed as a path consisting of two edges connecting different positions inside an attention head, and we remove paths inside attention heads following [Bakalova et al., 2025]. Removing all induction heads is thus done in two forward passes:
  • In all heads, replace V activation of a token in position P in the input string with zero when queried by a token in position P + 1 in the input string. Save K and V activations of all heads in position P + 1 in the input string.
  • In all heads, replace the K activation of a token in position P + 1 in the input string with the activation saved in previous step when queried by a token in position P in target string.
When removing anti-induction heads the second step is different: In all heads, replace V activation of a token in position P + 1 in input string with the activation saved in previous step when queried by a token in position P + 1 in target string.
@@
Figure 14: Difference in sum of attention scores per layer between fine-tuned and base models. The higher the difference, the more apmlified is this head in this layer during fine-tuning.
@@ Appendix F Compute Resources
We run all the prompting experiments using 7B/8B parameter models on a single H100 GPU and 32B/70B parameter models on 4xH100 GPUs with full precision without any quantization. Our fine-tuning experiments also use a single H100 GPU for all the experiments. All of our experiments should be reproducible with approximately 1500 H100 GPUh.

### v3

@@ C.4.2 Results: word-level vocabulary
To stress‑test copying while keeping text realistic, we construct a 1,500‑example corpus of synthetic Lorem Ipsum paragraphs. Generation starts from one base paragraph produced by the lorem Python library and then applies light, probabilistic perturbations to increase the non-determinism of the texts.
@@
  • Token budget. Sequences are truncated to 500 BPE tokens to stay within context limits while preserving paragraph‑level coherence.
@@
We identified successful prompts for copying naturalistic text as follows. We downloaded research papers put up on ArXiv in April 2025, and checked if the prompt results in perfect accuracy (not even a single mistake). Our motivation was that ArXiv text includes challenging text with equations, and, due to recency, should not have appeared in the LLM training data. More specifically, we segregated the sections of the papers into paragraphs, making sure the total length of the paragraph to be copied has close to 500 tokens (while ensuring that we do not end the paragraph in the middle of a sentence). We then pre-validated our prompts on a dataset containing $500$ samples each containing $500$ tokens, and even the smaller models we test – Llama3.1-8B and Qwen2.5-7B perfectly copied the given text (not a single mistake). We proceeded with our 3 minimal prompts, shown in Table 9, which achieved perfect copying accuracy. Each of our prompts ends with an unresolved <start> token, prompting the model to emit a verbatim repetition of the preceding paragraph.
Ratio of Ambiguous vs. Unambiguous Tokens. Table 7 reports, for each model, the average number of ambiguous (tokens with multiple possible continuations) and unambiguous (tokens with a single continuation) tokens per paragraph, aggregated across all random seeds. While different model families employ distinct tokenizers, so their absolute counts vary, yet the relative balance between ambiguous and unambiguous tokens is remarkably consistent. Each paragraph contains roughly 500 tokens, but repetition of sentences (to make the dataset appropriate for Non-unique copying) reduces the number of unique token types to only about 70. However, amongst those, both token types occur frequently in our data and the fact that copying “glitch” can be traced exclusively to unambiguous tokens is remarkable.
Table 7: Average per‐paragraph bigram counts by model
@@
Table 8: Model Performance at Longer (2k/3k/5k) Input Lengths
Unambiguous Ambiguous
Model 2k 3k 5k 2k 3k 5k
llama3-70B-Instruct 0.99 0.99 0.99 0.96 0.94 0.76
llama3-8B-Instruct 0.96 0.98 0.99 0.89 0.69 0.44
qwen2.5-32B-Instruct 1.00 0.99 0.99 0.95 0.94 0.93
qwen2.5-7B-Instruct 0.99 0.99 0.99 0.75 0.43 0.19
@@
echo‑tag RULE : Paragraphs are repeated **exactly** as is, whenever an Echo is seen.
@@
p‑tag‑the_end Please repeat the following paragraph exactly as it is, without any changes:
@@
C.5.1 Additional results for longer lengths
We conducted additional experiments with the same Lorem-Ipsum setup described above for a single seed at even longer input lengths (up to 5000 tokens) and report the results in Table 8. We observe the same effects we see at lengths of 500, that is, a lower success rate for ambiguous bigrams than for unambiguous ones. These results on longer inputs align perfectly with and amplify our findings in Section 4.2.
C.6 Prompting: Git Commit Revert vs Cherry-Pick (Section 4.2)
@@ C.5.1 Additional results for longer lengths
[…]
@@
We adopt exactly two natural‑language templates; one for the cherry‑pick orientation, one for the revert orientation. Each prompt describes the task in plain English, includes up to three few‑shot examples (hand‑crafted, $\ell=6$), shows the new history under a === NEW HISTORY === delimiter, and terminates with === ANSWER ===, where the model must emit the ordered commit list only. An example is given below.
@@
Because correct output demands global re‑ordering rather than token‑level copy, we found no benefit in adding separator toggles or mathematical re‑statements.
@@
Appendix D Fine-Tuning Details (Section 4.3)
@@ D.4 Results
We provide the training loss curves for each fine-tuning task in Figure 25.
Figure 25: Averaged training loss curve for GPT-2-XL model across all our tasks. For all the tasks, we achieve 0 training loss. Yet, the generalization behavior varies in line with theoretical predictions (Figure 6).
Appendix E Further Details on Mechanistic Interpretability Analysis (Section 4.4)
@@
To remove an induction head, we use path patching methodology, widely adopted in the literature [Hanna et al., 2023, Wang et al., 2022]. An induction head can be viewed as a path consisting of two edges connecting different positions inside an attention head: one edge connecting two adjacent positions in the input string, and one edge connecting the second of these with a position in the target string. We remove paths inside attention heads following Bakalova et al. [2025]. Removing all induction heads is thus done in two forward passes:
  • In all heads, replace the V activation of a token at position $P$ in the input string with zero whenever it is queried by the token in position $P+1$ in the input string. In the same forward pass, for each head, save K and V activations at position $P+1$ in the input string. Intuitively, this intervention prepares $K$ and $V$ activations at each position that carry no information about the preceding token.
  • In a second forward pass, in each head, replace the K activation of the token in position $P+1$ in the input string with the activation saved in the previous step whenever it is queried by a token in position $P$ in the target string. Intuitively, this intervention ensures that a head cannot attend to a position on the basis of information about the token immediately preceding the key position. That is precisely the behavior of induction heads [Olsson et al., 2022], i.e., induction heads are removed by this intervention.
When removing anti-induction heads, the first step is the same, but the second step is different: In each heads, we replace the V activation of the token in position $P+1$ in the input string with the activation saved in the first step whenever it is queried by a token in position $P+1$ in the target string. Intuitively, this intervention ensures that an attention head attending to a prior occurrence of the same token cannot retrieve the immediately preceding token.
The above description applies to Unique Forward Copy; in Unique Backward Copy, the position indices in the target string are appropriately reversed.
@@
Figure 26: Difference in sum of attention scores per layer between fine-tuned and base models. The higher the difference, the more amplified is this head in this layer during fine-tuning.
@@ Appendix F Compute Resources
We run all the prompting experiments using 4B/7B/8B parameter models on a single H100 GPU and 14B/32B/70B parameter models on 4xH100 GPUs with full precision without any quantization. Our fine-tuning experiments also use a single H100 GPU for all the experiments. All of our experiments should be reproducible with approximately 2500 H100 GPUh.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction assert that (1) fundamental length-generalization limits of small transformers may persist in large‐scale pretrained LLMs (2) pretrained models exhibit a notable induction-vs-anti-induction asymmetry; (3) this asymmetry can be traced to the relative strength of underlying attention circuits; and (4) targeted fine-tuning can rebalance these circuits and mitigate architectural limits, but does not entirely eliminate them. Through a mixture of synthetic as well as real world experiments in section 4.1, 4.2 we show point 2. Through our finetuning experiments in section 4.3 we justify point 4. We justify point 3 with experiments and discussion in section 4.4. Finally in totality the cumulation of all of our experiments and theoretical results justifies point 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss limitations of our work in Section 5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We base our theory on the extension of Huang et al. [2025]. In the main body of the paper we provide informal theorems and intuitions behind our ideas (Section 3). Detailed definitions, proofs, and remarks follow in Appendix A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide technical details of our tasks, in-context and fine-tuning experiments in Section 4. Specifically, we address the setup of the copying tasks in subsection 4.2 and fine-tuning setup in subsection 4.3. Additional details on prompting and fine-tuning are provided in the Appendix C and D respectively.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We attach the code and data used all across our experiments in the supplementary section. We also have provided sufficient details all throughout the paper in each section to help the readers reproduce any result.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have provided exact hyperparameters in the appendix, along with all the relevant training and test details for all experiments considering training in the appendix in sections D, B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We provide exact error bars in our experiments whenever relevant. We use at least 3 seeds for each experiment of ours to confirm their validity.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer:[Yes]
  39. Justification: We provide details of the compute resources used for all our experiments in our Appendix in Section F.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: There is no deviation from the NeurIPS Code of Ethics guidelines mentioned. We conform to all the guidelines.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Our paper as such connects theory of Transformers and contextualizes them for LLMs. Although we do not foresee any immediate negative positive societal impacts of our work. Long term, we hope that our insights that can be used to make better LLMs that when deployed in some critical domains can avoid the kind of errors we expose in LLMs
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A] .
  54. Justification: We do not foresee any misuse of the new benchmarks we release to expose some limitations in LLMs, and thus do not see the need to provide any safeguards related to the same.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the creators of models and code we use.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We contribute some new small benchmarks related to better understand limitations in LLMs, and attach the benchmarks, along with how to re-generate them in the supplementary material. We also attach all the code used to run our experiments in the supplementary section.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A] .
  69. Justification: We conduct no such experiments.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A] .
  74. Justification: We conduct no such experiments.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Our paper is about studying LLMs, and hence all of our experiments involved running LLMs to understand their behavior. Additionally we used LLMs to assist in refining our writing throughout the paper. LLMs were not used in any other part of the paper, including our theoretical results, coming up with our experimental design, creating our datasets and writing code.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.